OpenStudy (anonymous):

What is the value of the expression –7 – 5 + 4? A. –16 B. –6 C. –8 D. –12

OpenStudy (youngstudier):

What do you think?

OpenStudy (adll23):

The Value of -7 - 5 + 4 is: -8. Solution: Simplify the following: -7-5+4 Group the negative terms in -7-5+4 together and factor out the minus sign. -7-5+4 = 4-(7+5): 4-(7+5) Evaluate 7+5. 7+5 = 12: 4-12 Factor out a negative sign in 4-12 in order [ to subtract. 4-12 = -(12-4): -(12-4) Subtract 4 from 12. | 0 | 12 | 1 | 2 - | | 4 | 0 | 8: Answer: | | -8

OpenStudy (haseeb96):

first -7-5 = -12 -12 +4 = -8 this is your correct answer